### v1

@@ V Conclusions And Future Work
This paper developed the first methods that accurately identify/eliminate just the problematic attacked sensor data (keep the rest) presented to a sequence estimation/regression algorithm under a powerful attack model based on known/observed attacks. The developed approaches were shown to protect a sequence estimation/regression algorithm designed for unattacked sensor data by allowing such an algorithm to operate using the deemed to be unattacked data. The developed approaches employ only unattacked training data to mitigate all attacks allowed under the powerful considered attack model where the attacked sensors and attacks can change for each time step, the attacker has complete control of after-attack sensor data, and the protection system has no prior knowledge of how many sensors are attacked and which sensors are more likely to be attacked. Such powerful sensor attacks have not been mitigated elsewhere in the literature.
@@
We first proposed a simple protection approach for attacks not endowed with knowledge of the details of our protection approach, but later we proposed additional processing for attackers with knowledge. This allows a lower complexity approach to be employed if the attacker does not have detailed knowledge of the protection approach. Experimental results were provided which demonstrated good performance for both the simple and additional processing. Our simple method is shown to achieve estimation performance which is indistinguishable (to two decimal places) when compared to that of the GE (which knows which sensors were attacked) in cases tested, which assume the attacker does not have knowledge of the protection scheme. For cases where the attacker has knowledge of the protection approach, numerical results show that the additional processing can be configured, by proper parameter choice, so that the worst-case degradation under a large number of sensors attacked can be made significantly smaller than the worst-case degradation under the simple processing, and close to GE, for the same number of attacked sensors with just a slight degradation under no attacks. Guarding against the worst-case is extremely important and if this performance is acceptable, then performance will always be acceptable. Our protection approach is relatively low-complexity, as it does not employ high-complexity searches over all possible subsets of sensors.
@@
It would be of interest to further study methods to ensure good predictions where predictions are used in our approaches and maybe to monitor these predictions to make sure nothing has gone wrong. One issue of concern is to make sure problematic attacks do not leak through our protection to impact our predictions. Fortunately, we have found that this can be ensured even under very aggressive attacks if the APCC-ADDITIONAL approach employs $\alpha$ and $\beta$ which provide the needed level of protection. On the other hand, it would be of great interest to carefully study methods which might further enhance performance on these cases. This seems possible.
In closing, it should be noted that our approach can be considered one level of protection which can be combined with other levels to enhance overall protection. As our approach focuses on difficult attacks, it would be interesting to understand the gains of combining multiple approaches in this manner.

### v2

@@ 5 Conclusions And Future Work
This paper developed methods that accurately identify/eliminate just the problematic attacked sensor data (keep the rest) presented to a sequence estimation/regression algorithm for all possible attacks under our attack model. The developed approaches were shown to protect a sequence estimation/regression algorithm designed for unattacked sensor data by allowing such an algorithm to operate using the deemed to be unattacked data. The developed approaches employ only unattacked training data to mitigate all attacks allowed under our attack model where the attacked sensors and attacks can change for each time step, the attacker has complete control of after-attack sensor data, and the protection system has no prior knowledge of how many sensors are attacked and which sensors are more likely to be attacked.
@@
We first proposed a simple protection approach for attacks not endowed with knowledge of the details of our protection approach, but later we proposed additional processing for attackers with knowledge. This allows a lower complexity approach to be employed if the attacker does not have detailed knowledge of the protection approach. Experimental results were provided which demonstrated good performance for both the simple and additional processing. Our simple method is shown to achieve estimation performance which is indistinguishable (to two decimal places) when compared to that of the GE (which knows which sensors were attacked) in cases tested, which assume the attacker does not have knowledge of the protection scheme. For cases where the attacker has knowledge of the protection approach, numerical results show that the additional processing can be configured, by proper parameter choice, so that the worst-case degradation under a large number of sensors attacked can be made significantly smaller than the worst-case degradation under the simple processing, and close to GE, for the same number of attacked sensors with just a slight degradation under no attacks. Guarding against the worst-case is extremely important and if this performance is acceptable, then performance will always be acceptable.
@@
It would be of interest to further study methods to ensure good predictions where predictions are used in our approaches and maybe to monitor these predictions to make sure nothing has gone wrong. One issue of concern is to make sure problematic attacks do not leak through our protection to impact our predictions. Fortunately, we have found that this can be ensured even under very aggressive attacks if the APCC-ADDITIONAL approach employs $\alpha$ and $\beta$ which provide the needed level of protection. On the other hand, it would be of great interest to carefully study methods which might further enhance performance in these cases. This seems possible. Finally, it would be useful to study attacks on the training data, which we did not consider here.
In closing, it should be noted that our approach can be considered one level of protection which can be combined with other levels to enhance overall protection. It would be interesting to understand the gains of combining multiple approaches in this manner.